# AVA: Attentive VLM Agent for Mastering StarCraft II

### v3

@@ Abstract
We introduce Attentive VLM Agent (AVA), a multimodal StarCraft II agent that aligns artificial agent perception with the human gameplay experience. Traditional frameworks such as SMAC rely on abstract state representations that diverge significantly from human perception, limiting the ecological validity of agent behavior. Our agent addresses this limitation by incorporating RGB visual inputs and natural language observations that more closely simulate human cognitive processes during gameplay. The AVA architecture consists of three integrated components: (1) a vision-language model enhanced with specialized self-attention mechanisms for strategic unit targeting and battlefield assessment, (2) a retrieval-augmented generation system that leverages domain-specific StarCraft II knowledge to inform tactical decisions, and (3) a dynamic role-based task distribution system that enables coordinated multi-agent behavior. The experimental evaluation in our proposed AVACraft environment, which contains 21 multimodal StarCraft II scenarios, demonstrates that AVA powered by foundation models (specifically Qwen-VL and GPT-4o) can execute complex tactical maneuvers without explicit training, achieving comparable performance to traditional MARL methods that require substantial training iterations. This work establishes a foundation for developing human-aligned StarCraft II agents and advances the broader research agenda of multimodal game AI. Our implementation is available at https://github.com/camel-ai/VLM-Play-StarCraft2.
starcraft2, multi-agent, benchmark, environment
@@ 1 Introduction
StarCraft II has established itself as a cornerstone benchmark for artificial intelligence systems, particularly in the domain of multi-agent reinforcement learning (MARL). While environments such as SMAC and PySC2 have facilitated significant advances in AI capabilities, they primarily rely on abstract state representations that fundamentally differ from human perception and decision-making processes. This abstraction, while computationally efficient, creates a significant disconnect between how AI agents and humans interact with and understand the game environment.
Recent breakthroughs in Vision-Language Models (VLMs) have demonstrated remarkable capabilities in bridging the gap between computer vision and natural language understanding. These models excel at tasks requiring complex reasoning, strategic planning, and adaptation to novel scenarios without explicit training. However, their potential in real-time strategy games, particularly in environments requiring both tactical decision-making and multi-agent coordination, remains largely unexplored.
We present Attentive VLM Agent (AVA), a novel StarCraft II multi-modal agent that fundamentally reimagines how AI agents perceive and interact with the game world. To evaluate the performance of AVA, we propose a multimodal StarCraft II environment, AVACraft, which makes three key innovations: First, we redesign the observation space to align with human cognitive processes, incorporating both RGB visual input and natural language descriptions. This multimodal approach enables agents to process battlefield information in a manner analogous to human players, facilitating more intuitive strategic decision-making. Second, we develop a comprehensive action space that supports complex military maneuvers, including unit targeting, formation control, and ability usage. This design allows for fine-grained tactical control while maintaining accessibility for high-level strategic planning. Third, we introduce an integrated agent architecture in AVA that combines the reasoning capabilities of VLMs with specialized mechanisms for multi-agent coordination. We incorporate self-attention mechanisms for strategic unit targeting, retrieval-augmented generation (RAG) for leveraging domain-specific knowledge, and dynamic role assignment for task distribution.
@@
  • We propose an integrated VLM-based agent architecture (AVA) that combines self-attention mechanisms, retrieval-augmented generation, and dynamic role assignment for effective tactical decision-making.
@@
Our work establishes a foundation for developing more human-aligned StarCraft II agents and contributes to the broader goal of creating AI systems that can understand and interact with complex game environments in ways that more closely mirror human cognition.
@@ 2 Related Work
Vision-Language Model Benchmarks
In recent years, various VLM benchmarks have emerged that target different domains. VLMbench pioneered compositional designs for robotic manipulation, introducing structured evaluation of vision-language reasoning in physical tasks (Zheng et al., 2022). MultiMedEval (Royer et al., 2024) further advanced this direction by establishing a comprehensive framework for medical VLMs across multiple modalities, emphasizing standardized evaluation protocols. These benchmarks demonstrate the potential of VLMs in handling complex, domain-specific tasks, inspiring our approach to designing a VLM-based framework for StarCraft II that emphasizes human-like perception and decision-making processes.
General-Purpose Agents
The development of general-purpose agents has made significant strides in computer interaction tasks. CRADLE (Tan et al., 2024) established a unified framework using screenshots as input and keyboard/mouse actions as output, demonstrating the feasibility of vision-based control. WebVoyager (He et al., 2024) extended this approach through a sophisticated multimodal understanding of web interfaces, while ShowUI (Lin et al., 2024) introduced specialized techniques for GUI interaction, including UI-guided visual token selection and interleaved vision-language-action streaming. AppAgent (Zhang et al., 2023; Li et al., 2024a) further refined these concepts through simplified action spaces for smartphone control. These advances in general-purpose agents provide valuable insights for our work, particularly in designing intuitive interfaces and action spaces for StarCraft II environments.
@@
The evolution of StarCraft II AI systems has seen multiple paradigm shifts. AlphaStar (Vinyals et al., 2019) achieved superhuman performance through a combination of imitation learning and multi-agent reinforcement learning, followed by numerous improvements (Mathieu et al., 2021; star Contributors, 2021; Han et al., 2020; Huang et al., 2023). SMAC (Samvelyan et al., 2019; Ellis et al., 2023) provided a standardized framework for research but relied on abstract state representations that diverged from human perception. Recent works have begun exploring language models in StarCraft II, with LLM Play SC2 (Ma et al., 2024), LLM-PySC2 (Li et al., 2024b), and LLM-SMAC (Deng et al., 2024a) demonstrating the potential of language models in game understanding and decision-making. SMAC-Hard (Deng et al., 2024b) further extended these challenges by introducing more complex scenarios. Our work builds upon these foundations while introducing a novel approach that combines VLMs with attention mechanisms, creating a more human-aligned interaction paradigm that bridges the gap between abstract state representations and natural human gameplay.
@@ 3 AVACraft Environment Design
Figure 1: Observation Space of AVACraft environment.
Table 1: Comparison of StarCraft II environments. While SMAC simplifies the interface for MARL training and PySC2 provides comprehensive API access, AVACraft adapts PySC2 for multimodal decision-making through RGB observations and structured unit tracking.
SMAC PySC2 AVACraft
Base Environment PySC2 SC2 API PySC2
Visual Rep. None Feature layers (32x32x13) RGB screenshots
Unit Info Basic stats (health, coords) Detailed features (50x7) Tracked units with IDs
Battle Modes PVE with little abilities Full game modes PvP, PvE with abilities
Design Focus MARL Full game control Multimodal decision
Advantages Simple, stable Comprehensive API Visual processing Unit tracking Intuitive control
We formalize our AVACraft environment as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},P,R)$, where agents receive observations and take actions in discrete time steps.
@@ 3.1 State and Observation Space
The true environment state $s\in\mathcal{S}$ contains the complete game information. However, agents only receive partial observations $o\in\mathcal{O}$ consisting of:
$o_{t}=(I_{t},T_{t},U_{t}),$ (1)
where:
  • $I_{t}\in\mathbb{R}^{H\times W\times 3}$ is an RGB image observation of the current game state with configurable resolution $(H,W)$.
  • $T_{t}$ is a natural language description of the game state, including faction-specific unit status.
  • $U_{t}=\{u_{1},...,u_{n}\}$ represents the set of visible unit information, where each $u_{i}$ contains:
@@
    with $id_{i}$ being the unit identifier, $type_{i}$ the unit class, $pos_{i}\in\mathbb{R}^{2}$ the position, $attr_{i}$ the unit attributes, and $status_{i}$ the current status values.
@@ 3.2 Action Space
The action space $\mathcal{A}$ is defined as the union of three action types:
@@
where:
  • Attack actions $a_{attack}\in\mathcal{A}_{attack}$ are defined as ordered pairs $(i,j)$ representing unit $i$ attacking unit $j$
  • Move actions $a_{move}\in\mathcal{A}_{move}$ can be either:
    • Grid-based: $(i,x,y)$ where $i$ is the unit ID and $(x,y)\in\{1,...,10\}^{2}$
    • Directional: $(i,d)$ where $d\in\{\text{UP},\text{RIGHT},\allowbreak\text{DOWN},\text{LEFT}\}$
  • Ability actions $a_{ability}\in\mathcal{A}_{ability}$ are triples $(I,\allowbreak ability\_type,target)$ where $target$ can be a position, unit ID, or null depending on the ability type
@@ 3.3 Reward Structure and Episode Termination
where the episodes terminate under three conditions: Victory: All enemy units eliminated; Defeat: All allied units eliminated; Draw: Time limit (300 seconds) exceeded.
@@ 3.4 Battle Modes and Technical Features
Our environment supports both Player vs. Player (PvP) and Player vs. AI (PvE) modes, built on top of PySC2. We designed 12 specialized micromanagement scenarios covering essential tactical elements: multi-unit coordination, ability usage (e.g., Marine stutter-step, Stalker blink), advanced maneuvers (split, hit-and-run), and terrain utilization. The implementation features an automated unit annotation system and a grid system for precise spatial control, alongside observation wrappers that capture RGB screenshots and generate natural language descriptions.
Figure 2: VLM-Agent Architecture integrates AVA, a knowledge-enhanced decision system, and a dynamic role assignment framework.
@@ 4 Attentive VLM Agent Architecture
Our Attentive VLM Agent (AVA) architecture integrates three key components: a self-attention mechanism for strategic unit targeting, a knowledge-enhanced decision system, and a dynamic role assignment framework. This design enables effective battlefield analysis and coordinated unit control while maintaining alignment with human decision-making processes.
4.1 AVA Mechanism
We propose AVA, an integrated system that combines skill planning and tactical decision-making through vision-language models. The mechanism processes battlefield information through two key stages to emulate human-like attention and decision-making processes.
First, we introduce a VLM Planner that evaluates the battlefield situation and generates specific micro-management skill plans:
@@ 4.1 AVA Mechanism
where the planner outputs a structured skill plan: {mdframed}[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt] { “primary_skill”: { “name”: “Focus Fire”, “description”: “Concentrating damage on specific targets”, “steps”: [ “Select highest priority target”, “Command all units to attack the same target” ] }, “secondary_skills”: […] }
Based on the planner’s output, the system then performs strategic unit identification:
@@
where each annotation $a_{i}=(p_{i},c_{i},b_{i})$ consists of unit position $p_{i}$, unit class $c_{i}$, and bounding box $b_{i}$.
The strategic unit identification is then performed through a skill-aware natural language prompt:
@@
where:
  • $I$ is the current game screenshot
  • $T$ is the text state description
  • $H$ represents action history
  • $A$ is the set of unit annotations
  • $Q$ is the tactical analysis prompt generated based on the skill plan
  • $S$ is the current skill plan from VLM Planner
The VLM outputs its analysis in natural language, integrating both skill-based strategy and tactical considerations. For example, in a complex battlefield scenario with multiple Protoss units facing diverse Terran forces, the VLM analyzes the situation and identifies critical targets: {mdframed}[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt, roundcorner=10pt] Unit: Marine_1 (Tag: 7)
@@ 4.2 Knowledge Integration through RAG
To enhance the VLM’s tactical understanding with domain expertise, we implement a RAG system that operates on the priority units identified by AVA. Given the priority unit set $U_{\text{priority}}\subseteq A$, we formulate the knowledge retrieval and integration process as:
@@
where $D$ represents the tactical decision guidance generated by combining the retrieved knowledge with the current game state representation.
For example, given a priority Marine unit $u\in U_{\text{priority}}$ identified by AVA, the system retrieves structured knowledge $K(u)$ including combat statistics (DPS: 9.8 (+1.6)), counter relationships (Strong against: [Hydralisk, Immortal, Marauder]), and matchup-specific insights. The system then generates tactical guidance such as:
@@
This knowledge-grounded synthesis enables the agent to make tactically sound decisions based on both immediate battlefield conditions and established strategic principles.

### v5

@@ Abstract
We introduce Attentive VLM Agent (AVA), a multimodal StarCraft II agent that aligns artificial agent perception with the human gameplay experience. Traditional frameworks such as SMAC rely on abstract state representations that diverge significantly from human perception, limiting the ecological validity of agent behavior. Our agent addresses this limitation by incorporating RGB visual inputs and natural language observations that more closely simulate human cognitive processes during gameplay. The AVA architecture consists of three integrated components: (1) a vision-language model enhanced with specialized Multimodal Priority Inference mechanisms for strategic unit targeting and battlefield assessment, (2) a retrieval-augmented generation system that leverages domain-specific StarCraft II knowledge to inform tactical decisions, and (3) a dynamic role-based task distribution system that enables coordinated multi-agent behavior. The experimental evaluation in our proposed AVACraft environment, which contains 21 multimodal StarCraft II scenarios including both PvE and PvP settings, demonstrates that AVA powered by foundation models (specifically Qwen-VL and GPT-4o) can execute complex tactical maneuvers without explicit training, achieving comparable performance to traditional MARL methods that require substantial training iterations. Detailed case analyses further validate our approach, showing emergent tactical behaviors like focus-fire targeting and formation control that closely resemble professional human gameplay strategies. This work establishes a foundation for developing human-aligned StarCraft II agents and advances the broader research agenda of multimodal game AI. Our implementation is available at https://anonymous.4open.science/r/VLM-SMAC.
@@ 1 Introduction
StarCraft II has established itself as a cornerstone benchmark for artificial intelligence systems, particularly in the domain of multi-agent reinforcement learning (MARL). While environments such as SMAC and PySC2 have facilitated significant advances in AI capabilities, they primarily rely on abstract state representations that fundamentally differ from human perception and decision-making processes. This abstraction, while computationally efficient, creates a significant disconnect between how AI agents and humans interact with and understand the game environment. Recent breakthroughs in Vision-Language Models (VLMs) have demonstrated remarkable capabilities in bridging the gap between computer vision and natural language understanding. These models excel at tasks requiring complex reasoning, strategic planning, and adaptation to novel scenarios without explicit training. However, their potential in real-time strategy games, particularly in environments requiring both tactical decision-making and multi-agent coordination, remains largely unexplored.
We present Attentive VLM Agent (AVA), a novel StarCraft II multi-modal agent that fundamentally reimagines how AI agents perceive and interact with the game world. To evaluate the performance of AVA, we propose a multimodal StarCraft II environment, AVACraft, which makes three key innovations: First, we redesign the observation space to align with human cognitive processes, incorporating both RGB visual input and natural language descriptions. This multimodal approach enables agents to process battlefield information in a manner analogous to human players, facilitating more intuitive strategic decision-making. Second, we develop a comprehensive action space that supports complex military maneuvers, including unit targeting, formation control, and ability usage. This design allows for fine-grained tactical control while maintaining accessibility for high-level strategic planning. Third, we introduce an integrated agent architecture in AVA that combines the reasoning capabilities of VLMs with specialized mechanisms for multi-agent coordination. We incorporate Multimodal Priority Inference mechanisms for strategic unit targeting, retrieval-augmented generation (RAG) for leveraging domain-specific knowledge, and dynamic role assignment for task distribution.
@@
  • We propose an integrated VLM-based agent architecture (AVA) that combines Multimodal Priority Inference mechanisms, retrieval-augmented generation, and dynamic role assignment for effective tactical decision-making.
@@
Our work establishes a foundation for developing more human-aligned StarCraft II agents and contributes to the broader goal of creating AI systems that can understand and interact with complex game environments in ways that more closely mirror human cognition. In the supplementary material, we provide the code implementation and video demos of our agent decision-making.
Figure 1: Observation Space of AVACraft environment.
@@ 2 Related Work
VLM Benchmarks and Agents
Recent advancements in Vision-Language Models (VLMs) have led to the development of various benchmarks and agent frameworks. VLMbench (Zheng et al., 2022) pioneered compositional evaluation for robotic manipulation tasks, while MultiMedEval (Royer et al., 2024) established standardized protocols for medical applications. For enhancing visual understanding, Set-of-Mark (Yang et al., 2023) introduced region-based prompting techniques that significantly improve GPT-4V’s grounding capabilities. Comprehensive evaluation frameworks like MME (Fu et al., 2024a) and Video-MME (Fu et al., 2024b) have further advanced the field by assessing both static and sequential visual reasoning across diverse scenarios. Building on these visual foundations, agent frameworks have emerged to enable practical applications. CRADLE (Tan et al., 2024) established a unified screenshot-to-action pipeline, while WebVoyager (He et al., 2024) and ShowUI (Lin et al., 2024) developed sophisticated techniques for web interface navigation and GUI interaction. AppAgent (Zhang et al., 2023; Li et al., 2024a) further refined these concepts through optimized action spaces for mobile environments.
@@
StarCraft II has served as a critical benchmark for artificial intelligence, particularly for multi-agent systems. AlphaStar (Vinyals et al., 2019) marked a milestone by achieving superhuman performance through a combination of imitation learning and MARL, inspiring numerous architectural improvements (Mathieu et al., 2021; star Contributors, 2021; Han et al., 2020; Huang et al., 2023). For standardized evaluation, the StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al., 2019; Ellis et al., 2023) provided a widely-adopted framework, though it primarily relied on abstract state representations that diverge from human perception. Recent advancements in this space include SMACv2, which introduced procedurally generated scenarios requiring adaptive closed-loop policies, and SMAC-Hard (Deng et al., 2024b), which extended the challenge with more complex tactical scenarios. The integration of language models with StarCraft II has been explored in works such as LLM Play SC2 (Ma et al., 2024), LLM-PySC2 (Li et al., 2024b), and LLM-SMAC (Deng et al., 2024a), demonstrating the potential of language-based reasoning for strategic game understanding.
VLMs for Complex Gaming Environments
Despite progress in both fields, StarCraft II presents unique challenges for VLMs due to its requirements for real-time decision-making and multi-agent coordination. Our work bridges this gap by introducing AVA, which leverages VLMs’ visual understanding while aligning agent perception with human gameplay experience, establishing a foundation for human-aligned StarCraft II agents without relying on abstract state representations.
@@ 3 AVACraft Environment Design
Previous StarCraft II AI research environments present significant limitations for human-aligned agent development. SMAC and SMACv2 employ abstract feature representations that create a perception gap between AI and human players, modifying unit attributes and using ”cheat mode” mechanics that deviate from actual gameplay. While SMACv2’s randomization tests generalization capabilities, these scenarios often fail to reflect competitive StarCraft II compositions or tactics employed in professional gameplay. The details of these limitations is provided in Appendix B.
Our AVACraft environment addresses these issues by processing RGB visual input and natural language observations, enabling pretrained Vision-Language Models (VLMs) to make tactical decisions with explicit reasoning that aligns with human perception and gameplay mechanics. And the map details can be found in D. We formalize our AVACraft environment as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},P,R)$, where agents receive observations and take actions in discrete time steps.
@@ 3.1 State and Observation Space
The true environment state $s\in\mathcal{S}$ contains the complete game information. However, agents only receive partial observations $o\in\mathcal{O}$ consisting of $o_{t}=(I_{t},T_{t},U_{t})$, where $I_{t}\in\mathbb{R}^{H\times W\times 3}$ is an RGB image observation of the current game state with configurable resolution $(H,W)$, capturing the visual battlefield representation. $T_{t}$ is a natural language description of the game state, including faction-specific unit status, battlefield conditions, and relevant tactical information. $U_{t}=\{u_{1},...,u_{n}\}$ represents the set of visible unit information, where each $u_{i}$ contains:
@@
with $id_{i}$ being the unit identifier, $type_{i}$ the unit class, $pos_{i}\in\mathbb{R}^{2}$ the position coordinates, $attr_{i}$ the unit attributes (attack damage, armor, etc.), and $status_{i}$ the current status values (health, shields, energy, etc.).
Table 1: Comparison of StarCraft II environments. While SMAC and SMACv2 focus on reinforcement learning with abstract representations, AVACraft enables multimodal decision-making that aligns with human perception.
SMAC SMACv2 PySC2 AVACraft
Base Environment PySC2 PySC2 SC2 API PySC2
Visual Rep. None None Feature layers (32x32x13) RGB screenshots
Unit Info Basic stats (health, coords) Basic stats (health, coords) Detailed features (50x7) Tracked units with IDs
Battle Modes PVE with limited abilities Procedural scenarios Full game modes PvP, PvE with abilities
Design Focus MARL Generalization Full game control Multimodal decision
Advantages Simple, stable Randomization Generalization Comprehensive API Human-aligned perception Intuitive control
@@ 3.2 Action Space
The action space $\mathcal{A}$ is defined as the union of three distinct action types:
@@
where attack actions $a_{attack}\in\mathcal{A}_{attack}$ are defined as ordered pairs $(i,j)$ representing unit $i$ attacking unit $j$, enabling precise targeting decisions. Move actions $a_{move}\in\mathcal{A}_{move}$ can be formulated in two complementary ways:
  • Grid-based: $(i,x,y)$ where $i$ is the unit ID and $(x,y)\in\{1,...,10\}^{2}$ represents a discrete position in the environment’s spatial grid.
  • Directional: $(i,d)$ where $d\in\{\text{UP},\text{RIGHT},\allowbreak\text{DOWN},\text{LEFT}\}$ facilitates relative movement commands.
Ability actions $a_{ability}\in\mathcal{A}_{ability}$ are triples $(i,\allowbreak ability\_type,target)$ where $target$ can be a position, unit ID, or null depending on the ability type, supporting unit-specific tactical capabilities.
@@ 3.3 Reward Structure and Episode Termination
where the episodes terminate under three conditions: Victory: All enemy units eliminated; Defeat: All allied units eliminated; Draw: Time limit (300 seconds) exceeded. This approach provides clear performance signals while allowing flexible tactical execution.
@@ 3.4 Battle Modes and Technical Features
Our environment supports both Player vs. Player (PvP) and Player vs. AI (PvE) modes, built on top of PySC2. We designed 12 specialized micromanagement scenarios covering essential tactical elements: multi-unit coordination, ability usage (e.g., Marine stutter-step, Stalker blink), advanced maneuvers (split, hit-and-run), and terrain utilization. The implementation features an automated unit annotation system and a grid system for precise spatial control, alongside observation wrappers that capture RGB screenshots and generate natural language descriptions of battlefield conditions.
Figure 2: VLM-Agent Architecture integrates Multimodal Priority Inference mechanisms, a knowledge-enhanced decision system, and a dynamic role assignment framework.
@@ 4 Attentive VLM Agent Architecture
Our Attentive VLM Agent (AVA) architecture integrates three key components: a Multimodal Priority Inference mechanism for strategic unit targeting, a knowledge-enhanced decision system, and a dynamic role assignment framework. This architecture is designed to emulate human-like decision-making processes in complex StarCraft II scenarios.
4.1 Multimodal Priority Inference Mechanism
We propose a comprehensive system that combines structured skill planning and tactical decision-making through vision-language models. The mechanism processes battlefield information through two key stages to identify and prioritize strategic elements: First, we implement a VLM Planner that evaluates the battlefield situation and generates specific micro-management skill plans:
@@ 4.1 Multimodal Priority Inference Mechanism
where the planner outputs a structured skill plan with primary and secondary tactical objectives: {mdframed}[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt] { “primary_skill”: { “name”: “Focus Fire”, “description”: “Concentrating damage on specific targets”, “steps”: [ “Select highest priority target”, “Command all units to attack the same target” ] }, “secondary_skills”: […] }
Based on the planner’s output, the system performs precise unit identification and classification:
@@
where each annotation $a_{i}=(p_{i},c_{i},b_{i})$ consists of unit position $p_{i}$, unit class $c_{i}$, and bounding box $b_{i}$ for accurate spatial localization.
The critical Multimodal Priority Inference process then integrates visual features with tactical objectives through skill-aware natural language prompting:
@@
where $I$ is the current game screenshot capturing the visual battlefield state, $T$ is the text state description providing situational context, $H$ represents action history for temporal reasoning, $A$ is the set of unit annotations from the detection process, $Q$ is the tactical analysis prompt generated based on the skill plan, and $S$ is the current skill plan from the VLM Planner. The VLM outputs its analysis in structured natural language, integrating battlefield assessment with tactical prioritization. For example, in a complex scenario with multiple Protoss units facing diverse Terran forces, the system generates prioritized target assessments: {mdframed}[backgroundcolor=verylightgray, linecolor=black, linewidth=1pt, roundcorner=10pt] Unit: Marine_1 (Tag: 7)
@@ 4.2 Knowledge Integration through RAG
To enhance tactical decision-making with domain expertise, we implement a Retrieval-Augmented Generation (RAG) system that operates on the priority units identified through Multimodal Priority Inference. Given the priority unit set $U_{\text{priority}}\subseteq A$, we formulate the knowledge retrieval and integration process as:
@@
where $D$ represents the tactical decision guidance generated by combining the retrieved knowledge with the current game state representation. For example, given a priority Marine unit $u\in U_{\text{priority}}$ identified through our inference mechanism, the system retrieves structured knowledge $K(u)$ including combat statistics (DPS: 9.8 (+1.6)), counter relationships (Strong against: [Hydralisk, Immortal, Marauder]), and matchup-specific insights. The system then generates tactical guidance such as: